# Vacuum amplitudes and time-like causal unitary in the loop-tree duality

### v1

@@ III Local UV renormalisation and mass renormalisation scheme
and then we expand the phase-space residues ${\cal A}_{\rm D}^{(3,f)}(456)$ in Eq. (19) and Eq. (20) for $\rho\to\infty$. This expansion provides the most singular UV behaviour from which the UV counterterm is constructed. The scale $\mu_{\rm UV}$ is interpreted as the renormalisation scale [18] in the sense that the UV counterterm suppresses all the energy modes of ${\cal A}_{\rm D}^{(3,f)}(456)$ above $\mu_{\rm UV}$, and locally matches its UV behaviour at very high energies. The UV counterterm renders the phase-space residue UV finite. Yet, it is also necessary to subtract subleading terms to deliver the expected results in, e.g., the ${\overline{\rm MS}}$ renormalisation scheme, after setting the space-time dimensions to $d=4$ in the integrand.
The UV counterterm for the Higgs boson decay reads
${\cal A}_{{\rm UV}}^{(3,H)}(456)=\frac{g_{H}^{(1)}}{x_{45}}\bigg{[}\Delta Z_{H}^{({\rm UV})}\,\overline{|{\cal M}_{H\to q\bar{q}}^{(0)}|^{2}}$
@@
The coefficients $c_{H}^{({\rm UV})}=d$ and $c_{\gamma}^{({\rm UV})}=(d-2)/2$ were defined in Ref. [24] and $\lambda_{{\rm UV}}=\sqrt{\bm{\ell}_{1}^{2}+\mu_{{\rm UV}}^{2}-\imath 0}$. The terms proportional to $\mu_{\rm UV}^{2}$ in Eq. (26) are UV subleading contribution that determine the renormalisation scheme. The function
${\bf\Delta}_{H}^{({\rm UV})}=\frac{c_{\gamma}^{({\rm UV})}}{s\lambda_{{\rm UV}}^{3}}\left(-\frac{3(\bm{\ell}_{1}\cdot\bm{\ell}_{2})^{2}}{\lambda_{{\rm UV}}^{2}}+\bm{\ell}_{2}^{2}+2\bm{\ell}_{1}\cdot\bm{\ell}_{2}\right)$
×$\times\overline{|{\cal M}_{H\to q\bar{q}}^{(0)}|^{2}}~{},$ (27)
@@
$\int d\Phi_{\bm{\ell}_{1}}{\cal A}_{{\rm UV}}^{(3,H)}(456)\widetilde{\Delta}_{45\bar{6}}=\frac{\pi}{s\beta}\,g_{H}^{(1)}\,\widetilde{S}_{\epsilon}\bigg{[}\frac{\mu^{2\epsilon}}{\mu_{\rm UV}^{2\epsilon}}$
×$\times\frac{3}{\epsilon}\overline{|{\cal M}_{H\to q\bar{q}}^{(0)}|^{2}}-\frac{\mu^{2\epsilon}}{m^{2\epsilon}}\left(\frac{3}{\epsilon}+4\right)8m^{2}(1+\beta^{2})\bigg{]}~{},$ (28)
@@
The vacuum amplitude correctly accounts for the derivative of the quark and antiquark selfenergies through the residues of the corresponding squared Feynman propagators. However, it also generates a term proportional to the square of the mass. This term would not be considered in the state-of-the-art approach because the renormalised selfenergy is fixed to vanish on shell. The term $\Delta Z_{m}^{({\rm UV})}$ in Eq. (26) restores the mass renormalisation in the on-shell scheme when the renormalisation scale is identified with the quark mass itself, which introduces an interesting physical interpretation. Moreover, the mass renormalisation in other schemes, e.g., the ${\overline{\rm MS}}$ scheme, would be achieved by modifying the subleading UV factors of $\Delta Z_{m}^{({\rm UV})}$.
In the case of an off-shell photon, the UV counterterm reads
${\cal A}_{{\rm UV}}^{(3,\gamma^{*})}(456)=\frac{g_{\gamma^{*}}^{(1)}}{x_{45}}\bigg{[}\Delta Z_{m}^{({\rm UV})}(-8m^{2})\left(1-\frac{\beta^{2}}{d-2}\right)$
$+{\bf\Delta}_{\gamma^{*}}^{({\rm UV})}\bigg{]}~{},$ (30)
@@
${\bf\Delta}_{\gamma^{*}}^{({\rm UV})}=\frac{c_{\gamma}^{({\rm UV})}}{s\lambda_{{\rm UV}}^{3}}\left(-\frac{3(\bm{\ell}_{1}\cdot\bm{\ell}_{2})^{2}}{\lambda_{{\rm UV}}^{2}}+\bm{\ell}_{2}^{2}+2\bm{\ell}_{1}\cdot\bm{\ell}_{2}\right)$
×$\times\overline{|{\cal M}_{\gamma^{*}\to q\bar{q}}^{(0)}|^{2}}+\frac{3(2\bm{\ell}_{1}\cdot\bm{\ell}_{2})^{2}}{\lambda_{{\rm UV}}^{2}}+\left(d-4+\frac{3\mu_{\rm UV}^{2}}{\lambda_{\rm UV}^{2}}-\beta^{2}\right)s~{},$
@@
×$\times\frac{\mu^{2\epsilon}}{m^{2\epsilon}}\left(\frac{3}{\epsilon}+4\right)(-8m^{2})\left(1-\frac{\beta^{2}}{d-2}\right)~{}.$ (32)
This result is consistent with the fact that conserved or partially conserved currents, such as the vector and axial currents, do not get renormalised. In other words, the UV counterterm in Eq. (30) renormalises the quark mass in the on-shell scheme, exactly as Eq. (25) does, but does not renormalise the wave function.
@@
IV A toy decay rate at NNLO
Figure 2: Four-loop vacuum diagram generating the decay rate of a massive scalar particle at NNLO. The gray dashed lines represent phase-space residues.
@@ IV A toy decay rate at NNLO
$+{\cal A}_{\rm D}^{(4,\Phi,{\rm R})}(1356)\,\widetilde{\Delta}_{135\bar{6}}+{\cal A}_{\rm D}^{(4,\Phi)}(35678)\,\widetilde{\Delta}_{35\bar{6}78}\Big{)}$
[…]
@@
${\cal A}_{\rm D}^{(4,\Phi)}(456)=\frac{g_{\Phi}^{(2)}s^{3}}{x_{1234578}}$
×$\times\Bigg{[}\frac{1}{\lambda_{13\bar{4}}}\left(L^{23\bar{4}\bar{5}}_{3\bar{4}78,2\bar{5}78}+L^{2578}_{125,3\bar{4}78}+L^{178}_{125,2\bar{5}78}\right)$
$+\frac{1}{\lambda_{134}}\left(L^{2578}_{2345,178}+L^{3478}_{2345,2\bar{5}78}\right)$
$+\frac{1}{\lambda_{12\bar{5}}}\left(L^{134}_{2\bar{5}78,178}+L^{23\bar{4}\bar{5}}_{2\bar{5}78,178}\right)+\frac{1}{\lambda_{125}}L^{2345}_{2578,178}$
$+\frac{1}{\lambda_{178}}\left(L^{3478}_{2345,2\bar{5}78}+L^{3\bar{4}78}_{23\bar{4}\bar{5},2578}\right)\Bigg{]}~{},$ (36)
@@
${\cal A}_{\rm D}^{(4,\Phi)}(1356)=\frac{g_{\Phi}^{(2)}s^{3}}{x_{13578}}\left(\frac{1}{\lambda_{13\bar{4}}\lambda_{134}\lambda_{1\bar{2}5}\lambda_{125}}\right)$
×$\qquad\times\left(\frac{1}{\lambda_{\bar{1}78}}+\frac{1}{\lambda_{178}}\right)~{},$ (38)
@@
${\cal A}_{\rm D}^{(4,\Phi)}(35678)=-\frac{g_{\Phi}^{(2)}s^{3}}{x_{3578}}$
×$\qquad\times\left(\frac{1}{\lambda_{3\bar{4}78}\lambda_{3478}\lambda_{\bar{2}578}\lambda_{2578}\lambda_{\bar{1}78}\lambda_{178}}\right)~{}.$ (39)
@@
$\lim_{\lambda_{13\bar{4}}\to 0}\left({\cal A}_{\rm D}^{(4,\Phi)}(456)\widetilde{\Delta}_{45\bar{6}}+{\cal A}_{\rm D}^{(4,\Phi)}(1356)\widetilde{\Delta}_{135\bar{6}}\right)$
$\qquad={\cal O}(\lambda_{13\bar{4}}^{0})~{},$
$\lim_{\lambda_{\bar{1}78}\to 0}\left({\cal A}_{\rm D}^{(4,\Phi)}(1356)\widetilde{\Delta}_{135\bar{6}}+{\cal A}_{\rm D}^{(4,\Phi)}(35678)\widetilde{\Delta}_{35\bar{6}78}\right)$
$\qquad={\cal O}(\lambda_{\bar{1}78}^{0})~{}.$ (40)
@@
$\lim_{\lambda_{3\bar{4}78}\to 0}\left({\cal A}_{\rm D}^{(4,\Phi)}(456)\widetilde{\Delta}_{45\bar{6}}+{\cal A}_{\rm D}^{(4,\Phi)}(1356)\widetilde{\Delta}_{135\bar{6}}\right.$
$\qquad\left.+{\cal A}_{\rm D}^{(4,\Phi)}(35678)\widetilde{\Delta}_{35\bar{6}78}\right)={\cal O}(\lambda_{3\bar{4}78}^{0})~{}.$ (41)
Finally, a unitary threshold singularity arises at $\lambda_{23\bar{4}\bar{5}}\to 0$ that matches locally between ${\cal A}_{\rm D}^{(4,\Phi)}(456)$ and ${\cal A}_{\rm D}^{(4,\Phi)}(236)$.
@@
${\cal A}_{{\rm UV}}^{(4,\Phi)}(456)=\frac{g_{\Phi}^{(2)}s}{4\lambda_{{\rm UV}}^{3}}{\cal A}_{\rm D}^{(3,\Phi)}(456)~{},$
${\cal A}_{{\rm UV}}^{(4,\Phi)}(1356)=\frac{g_{\Phi}^{(2)}s}{4\lambda_{{\rm UV}}^{3}}{\cal A}_{\rm D}^{(3,\Phi)}(1356)~{},$ (42)
@@
$d\Gamma^{(1)}_{\Phi\to\phi\phi}=\frac{d\Phi_{\bm{\ell}_{1}\bm{\ell}_{2}}}{2\sqrt{s}}\,\bigg{[}\Big{(}{\cal A}_{\rm D}^{(3,\Phi)}(456)\,\widetilde{\Delta}_{45\bar{6}}$
$\quad+{\cal A}_{\rm D}^{(3,\Phi)}(1356)\,\widetilde{\Delta}_{135\bar{6}}\Big{)}+(5\leftrightarrow 2,4\leftrightarrow 3)\bigg{]}~{}.$ (44)
V Numerical implementation
Figure 3: Numerical implementation of LTD causal unitary at NLO for the three decay processes $H\to q\bar{q}$, $\gamma^{*}\to q\bar{q}$ and $\Phi\to\phi\phi$ as a function of the final state mass.
With the expressions presented in the previous sections, we now introduce a numerical implementation directly in the three physical spacial dimensions, $d-1=3$, since IR singularities cancel locally among phase-space residues and the local UV counterterm properly accounts for the singular UV behaviour. Unitary threshold singularities also match locally in the sum of the loop phase-space residues rendering the integrand well behaved across thresholds. To this end, we must consider the constraints introduced by energy conservation, which are encoded by the functions $\widetilde{\Delta}_{i_{1}\cdots i_{n}\bar{6}}$ in Eq. (15), Eq. (18) and Eq. (35).

### v2

@@ 3 Local UV renormalisation and mass renormalisation scheme
This expansion provides the most singular UV behaviour from which the UV counterterm is constructed. The scale $\mu_{\rm UV}$ is interpreted as the renormalisation scale Hernandez-Pinto:2015ysa in the sense that the UV counterterm suppresses all the energy modes of ${\cal A}_{\rm D}^{(3,f)}(456)$ above $\mu_{\rm UV}$, and locally matches its UV behaviour at very high energies. The UV counterterm renders the phase-space residue UV finite. Yet, it is also necessary to subtract subleading terms to deliver the expected results in, e.g., the ${\overline{\rm MS}}$ renormalisation scheme, after setting the space-time dimensions to $d=4$ in the integrand. It is important to stress that we do not perform the UV expansion diagram by diagram, but consider the vacuum amplitude including all contributions. This approach is more efficient because it takes into account possible cancellations in the UV regions of the loop momenta and leads to more compact local UV counterterms.
The local UV counterterm for the Higgs boson decay reads
@@
The coefficients $c_{H}^{({\rm UV})}=d$ and $c_{\gamma}^{({\rm UV})}=(d-2)/2$ were defined in Ref. Driencourt-Mangin:2019aix and $\lambda_{{\rm UV}}=\sqrt{\bm{\ell}_{1}^{2}+\mu_{{\rm UV}}^{2}-\imath 0}$. The terms proportional to $\mu_{\rm UV}^{2}$ in Eq. (26) are UV subleading contribution that determine the renormalisation scheme. The function
${\bf\Delta}_{H}^{({\rm UV})}=\frac{c_{\gamma}^{({\rm UV})}}{s\lambda_{{\rm UV}}^{3}}\left(-\frac{3(\bm{\ell}_{1}\cdot\bm{\ell}_{2})^{2}}{\lambda_{{\rm UV}}^{2}}+\bm{\ell}_{2}^{2}+2\bm{\ell}_{1}\cdot\bm{\ell}_{2}\right)\overline{|{\cal M}_{H\to q\bar{q}}^{(0)}|^{2}}~{},$ (27)
@@
$\int d\Phi_{\bm{\ell}_{1}}{\cal A}_{{\rm UV}}^{(3,H)}(456)\widetilde{\Delta}_{45\bar{6}}=\frac{\pi}{s\beta}\,g_{H}^{(1)}\,\widetilde{S}_{\epsilon}\bigg{[}\frac{\mu^{2\epsilon}}{\mu_{\rm UV}^{2\epsilon}}\frac{3}{\epsilon}\overline{|{\cal M}_{H\to q\bar{q}}^{(0)}|^{2}}$
$\quad-\frac{\mu^{2\epsilon}}{m^{2\epsilon}}\left(\frac{3}{\epsilon}+4\right)8m^{2}(1+\beta^{2})\bigg{]}~{},$ (28)
@@
The vacuum amplitude correctly accounts for the derivative of the quark and antiquark selfenergies through the residues of the corresponding squared Feynman propagators. However, it also generates a term proportional to the square of the mass. This term would not be considered in the state-of-the-art approach because the renormalised selfenergy is fixed to vanish on shell. The term $\Delta Z_{m}^{({\rm UV})}$ in Eq. (26) restores the mass renormalisation in the on-shell scheme when the renormalisation scale is identified with the quark mass itself. The requirement that $\mu_{\rm UV}=m$ in $\Delta Z_{m}^{({\rm UV})}$ also suggests an interesting physical interpretation.
In the case of an off-shell photon, the local UV counterterm reads
${\cal A}_{{\rm UV}}^{(3,\gamma^{*})}(456)=\frac{g_{\gamma^{*}}^{(1)}}{x_{45}}\bigg{[}\Delta Z_{m}^{({\rm UV})}(-8m^{2})\left(1-\frac{\beta^{2}}{d-2}\right)+{\bf\Delta}_{\gamma^{*}}^{({\rm UV})}\bigg{]}~{},$ (30)
@@
${\bf\Delta}_{\gamma^{*}}^{({\rm UV})}=\frac{c_{\gamma}^{({\rm UV})}}{s\lambda_{{\rm UV}}^{3}}\left(-\frac{3(\bm{\ell}_{1}\cdot\bm{\ell}_{2})^{2}}{\lambda_{{\rm UV}}^{2}}+\bm{\ell}_{2}^{2}+2\bm{\ell}_{1}\cdot\bm{\ell}_{2}\right)\overline{|{\cal M}_{\gamma^{*}\to q\bar{q}}^{(0)}|^{2}}$
$\quad+\frac{3(2\bm{\ell}_{1}\cdot\bm{\ell}_{2})^{2}}{\lambda_{{\rm UV}}^{2}}+\left(d-4+\frac{3\mu_{\rm UV}^{2}}{\lambda_{\rm UV}^{2}}-\beta^{2}\right)s~{},$ (31)
@@
This result is consistent with the fact that conserved or partially conserved currents, such as the vector and axial currents, do not get renormalised. In other words, the UV counterterm in Eq. (30) renormalises the quark mass in the on-shell scheme, exactly as Eq. (25) does, but does not renormalise the wave function because the term proportional to the squared LO is absent. Note that the vertex and selfenergy diagrams contribute with opposite signs in the UV, and therefore by considering a unitary approach there is a manifest cancellation between them without the need to subtract their UV behaviour separately. Note also that for massless quarks the local UV counterterms in Eq. (25) and Eq. (30) are well defined, in particular, the term containing $\Delta Z_{m}^{({\rm UV})}$ is proportional to the quark mass and thus has an smooth massless limit. The integrated UV counterterms in Eq. (28) and Eq. (32) behave as $m^{2}\mathrm{log}\left(m^{2}\right)$, while the customary wave function and mass renormalisation constants Sborlini:2016hat behave as $\mathrm{log}\left(m^{2}\right)$ and are therefore ill-defined in the massless limit. This is another advantage of the LTD causal unitary approach over other implementations.
@@
Beyond the one loop, the construction of the local UV counterterm, although more subtle because it requires subtracting the UV behaviour of the phase-space residues in different combinations of UV configurations of the loop momenta, is algorithmically well defined in terms of the expansions defined by Eq. (24), as shown in detail in Ref. Rios-Sanchez:2024xtv .
4 A toy decay rate at NNLO
Figure 2: Four-loop vacuum diagram contributing to the decay rate of a massive scalar particle at NNLO. The gray dashed lines represent phase-space residues.
@@ 4 A toy decay rate at NNLO
$\quad+{\cal A}_{\rm D}^{(4,\Phi,{\rm R})}(1356)\,\widetilde{\Delta}_{135\bar{6}}+{\cal A}_{\rm D}^{(4,\Phi)}(35678)\,\widetilde{\Delta}_{35\bar{6}78}\Big{)}+(5\leftrightarrow 2,4\leftrightarrow 3)\bigg{]}~{}.$ (35)
@@
${\cal A}_{\rm D}^{(4,\Phi)}(456)=\frac{g_{\Phi}^{(2)}s^{3}}{x_{1234578}}\Bigg{[}\frac{1}{\lambda_{13\bar{4}}}\left(L^{23\bar{4}\bar{5}}_{3\bar{4}78,2\bar{5}78}+L^{2578}_{125,3\bar{4}78}+L^{178}_{125,2\bar{5}78}\right)$
$\quad+\frac{1}{\lambda_{134}}\left(L^{2578}_{2345,178}+L^{3478}_{2345,2\bar{5}78}\right)+\frac{1}{\lambda_{12\bar{5}}}\left(L^{134}_{2\bar{5}78,178}+L^{23\bar{4}\bar{5}}_{2\bar{5}78,178}\right)+\frac{1}{\lambda_{125}}L^{2345}_{2578,178}$
$\quad+\frac{1}{\lambda_{178}}\left(L^{3478}_{2345,2\bar{5}78}+L^{3\bar{4}78}_{23\bar{4}\bar{5},2578}\right)\Bigg{]}~{},$ (36)
@@
${\cal A}_{\rm D}^{(4,\Phi)}(1356)=\frac{g_{\Phi}^{(2)}s^{3}}{x_{13578}}\left(\frac{1}{\lambda_{13\bar{4}}\lambda_{134}\lambda_{1\bar{2}5}\lambda_{125}}\right)\left(\frac{1}{\lambda_{\bar{1}78}}+\frac{1}{\lambda_{178}}\right)~{},$ (38)
@@
${\cal A}_{\rm D}^{(4,\Phi)}(35678)=-\frac{g_{\Phi}^{(2)}s^{3}}{x_{3578}}\left(\frac{1}{\lambda_{3\bar{4}78}\lambda_{3478}\lambda_{\bar{2}578}\lambda_{2578}\lambda_{\bar{1}78}\lambda_{178}}\right)~{}.$ (39)
@@
$\lim_{\lambda_{13\bar{4}}\to 0}\left({\cal A}_{\rm D}^{(4,\Phi)}(456)\widetilde{\Delta}_{45\bar{6}}+{\cal A}_{\rm D}^{(4,\Phi)}(1356)\widetilde{\Delta}_{135\bar{6}}\right)={\cal O}(\lambda_{13\bar{4}}^{0})~{},$
$\lim_{\lambda_{\bar{1}78}\to 0}\left({\cal A}_{\rm D}^{(4,\Phi)}(1356)\widetilde{\Delta}_{135\bar{6}}+{\cal A}_{\rm D}^{(4,\Phi)}(35678)\widetilde{\Delta}_{35\bar{6}78}\right)={\cal O}(\lambda_{\bar{1}78}^{0})~{}.$ (40)
@@
$\lim_{\lambda_{3\bar{4}78}\to 0}\left({\cal A}_{\rm D}^{(4,\Phi)}(456)\widetilde{\Delta}_{45\bar{6}}+{\cal A}_{\rm D}^{(4,\Phi)}(1356)\widetilde{\Delta}_{135\bar{6}}+{\cal A}_{\rm D}^{(4,\Phi)}(35678)\widetilde{\Delta}_{35\bar{6}78}\right)={\cal O}(\lambda_{3\bar{4}78}^{0})~{}.$
Finally, a threshold singularity arises at $\lambda_{23\bar{4}\bar{5}}\to 0$ that matches locally between ${\cal A}_{\rm D}^{(4,\Phi)}(456)$ and ${\cal A}_{\rm D}^{(4,\Phi)}(236)$.
@@
${\cal A}_{{\rm UV}}^{(4,\Phi)}(456)=\frac{g_{\Phi}^{(2)}s}{4\lambda_{{\rm UV}}^{3}}{\cal A}_{\rm D}^{(3,\Phi)}(456)~{},\qquad{\cal A}_{{\rm UV}}^{(4,\Phi)}(1356)=\frac{g_{\Phi}^{(2)}s}{4\lambda_{{\rm UV}}^{3}}{\cal A}_{\rm D}^{(3,\Phi)}(1356)~{},$ (42)
@@
$d\Gamma^{(1)}_{\Phi\to\phi\phi}$ $=$ $\frac{d\Phi_{\bm{\ell}_{1}\bm{\ell}_{2}}}{2\sqrt{s}}\,\bigg{[}\Big{(}{\cal A}_{\rm D}^{(3,\Phi)}(456)\,\widetilde{\Delta}_{45\bar{6}}+{\cal A}_{\rm D}^{(3,\Phi)}(1356)\,\widetilde{\Delta}_{135\bar{6}}\Big{)}+(5\leftrightarrow 2,4\leftrightarrow 3)\bigg{]}~{}.$ (44)
5 Numerical implementation
Figure 3: Numerical implementation of LTD causal unitary at NLO for the three decay processes $H\to q\bar{q}$, $\gamma^{*}\to q\bar{q}$ and $\Phi\to\phi\phi$ as a function of the final state mass. The photon current is conserved and the scalar decay is UV finite. Consequently, they do not exhibit any dependence on the renormalisation scale $\mu_{\rm UV}$.
With the expressions presented in the previous sections, we now introduce a numerical implementation directly in the three physical spacial dimensions, $d-1=3$, since IR singularities cancel locally among phase-space residues and the local UV counterterm properly accounts for the singular UV behaviour. Threshold singularities also match locally in the sum of the loop phase-space residues rendering the integrand well behaved across thresholds. To this end, we must consider the constraints introduced by energy conservation, which are encoded by the functions $\widetilde{\Delta}_{i_{1}\cdots i_{n}\bar{6}}$ in Eq. (15), Eq. (18) and Eq. (35).
$2m/\sqrt{s}=$ $10^{-3}$ $0.2$ $0.4$ $0.6$ $0.8$
$\Phi\to\phi\phi$ 00 $0.001760$ $0.006509$ $0.016693$ $0.045894$
$0.000029(14)$ $0.001785(10)$ $0.006517(7)$ $0.016702(8)$ $0.045929(10)$
$\gamma^{*}\to q\bar{q}(g)$ $0.018998$ $0.021524$ $0.031273$ $0.054715$ $0.114001$
$0.018963(57)$ $0.021556(24)$ $0.031280(22)$ $0.054785(36)$ $0.114013(19)$
$H\to q\bar{q}(g)$ $0.054981$ $0.067702$ $0.095937$ $0.140174$ $0.220367$
$\mu_{\rm UV}=2m_{H}$ $0.055017(33)$ $0.067713(17)$ $0.095948(16)$ $0.140200(16)$ $0.220364(16)$
$H\to q\bar{q}(g)$ $0.107654$ $0.120374$ $0.148610$ $0.192847$ $0.273040$
$\mu_{\rm UV}=m_{H}$ $0.107642(45)$ $0.120361(24)$ $0.148629(23)$ $0.192858(29)$ $0.273013(24)$
$H\to q\bar{q}(g)$ $0.160327$ $0.173047$ $0.201283$ $0.245520$ $0.325713$
$\mu_{\rm UV}=m_{H}/2$ $0.160305(56)$ $0.173070(50)$ $0.201319(45)$ $0.245522(44)$ $0.325708(41)$
Table 1: Numerical implementation of LTD causal unitary at NLO for the three decay processes $H\to q\bar{q}$, $\gamma^{*}\to q\bar{q}$ and $\Phi\to\phi\phi$ as a function of the final state mass. The first row in each item is the analytical result in DREG from the expressions presented in Appendix B. The photon current is conserved and the scalar decay is UV finite. Consequently, they do not exhibit any dependence on the renormalisation scale $\mu_{\rm UV}$.